# Universal Few-Shot Spatial Control for Diffusion Models

### v1

@@ 5.2 Main Results
Fully-supervised ControlNet Zhang et al. (2023) 0.3598 0.5972 89.09 14.14 0.525 0.4824
Uni-ControlNet Zhao et al. (2023) 0.3378 0.5808 92.60 14.59 0.340 0.4695
Training-free Ctrl-X Lin et al. (2024) 0.2901 0.3002 98.10 19.38 0.005 0.1352
Few-shot Prompt Diffusion Wang et al. (2023) 0.2120 0.2887 98.81 20.34 0.012 0.2266
Uni-ControlNet + FT Zhao et al. (2023) 0.2222 0.2917 99.21 20.36 0.010 0.2687
Prompt Diffusion + FT Wang et al. (2023) 0.2773 0.4810 95.64 18.04 0.034 0.3548
@@
Fully-supervised ControlNet Zhang et al. (2023) 21.06 17.15 20.16 21.73 44.33 34.72
Uni-ControlNet Zhao et al. (2023) 16.79 16.76 20.35 19.59 47.13 36.77
Training-free Ctrl-X Wang et al. (2023) 29.83 30.18 30.32 30.35 56.60 45.88
Few-shot Prompt Diffusion Wang et al. (2023) 24.85 23.00 24.07 23.28 43.60 34.15
Uni-ControlNet Zhao et al. (2023) + FT 24.58 25.48 24.77 25.12 47.82 36.06
Prompt Diffusion Wang et al. (2023) + FT 20.57 19.86 24.56 24.30 45.94 36.83
@@
We report the 30-shot performance of our model with UNet backbone Rombach et al. (2022) in terms of two aspects: controllability (Table 1) and image quality (Table 2). Despite using only a few annotated examples, UFC achieves comparable image quality and over 90% of the controllability performance of fully supervised baselines across most conditions. We observe that dense condition maps benefit more from the matching mechanism, as fine-grained similarity modeling enables more effective control features. This is reflected in the smaller relative performance gap with fully supervised baselines on Densepose compared to Pose.
@@ 5.3 Analysis
To assess the effectiveness of our design in handling the large discrepancy across condition modalities, we conduct an ablation study with two variants of UFC on UNet backbone Rombach et al. (2022): (1) UFC w/o Matching and (2) UFC w/o Fine-tuning. Both variants are trained on the same dataset as our main method and evaluated on the COCO 2017 validation split. In (1), we remove the matching mechanism and encode the query condition directly using the condition encoder i.e., $\mathcal{G}_{\tau}(y_{\tau})$ while retaining task-specific parameters and meta-learning paradigm. At test time, we fine-tune $\theta_{\tau}$ and $\mathcal{Z}$ using the small support set. In (2), we disable adaptation entirely by sharing all parameters across tasks during training and evaluating directly on unseen tasks without any fine-tuning.
@@
Since UFC is applicable to various diffusion backbones, we employ the stronger pre-trained model based on DiT Esser et al. (2024) and compare it with the one based on UNet used in our main table. The quantitative results are reported in Table 4, showing that the DiT-based model with a stronger diffusion backbone consistently outperforms the UNet-based model in controllability across all tasks. The lower image quality compared to the UNet variant, which is also observed in Tan et al. (2024), can be due to the mismatch between the optimized image resolution of the pre-trained diffusion backbone with the testing resolution (1024 and 512). Figure 9 in Appendix D shows that UFC (DiT) qualitatively achieves more fine-grained controllability than the UNet counterpart. More qualitative examples using the DiT backbone are shown in Figure 10 in Appendix.
@@
To further validate the few-shot capability of our method in more challenging settings, we evaluate our method on novel spatial conditions on 3D structures, such as 3D meshes, wireframes, and point clouds extracted from iso3d dataset Ebert (2025). Importantly, these control tasks involve not only novel spatial conditions but also different output image distributions from meta-training datasets, simulating a more realistic generalization scenario to unseen control tasks.
@@ A.1 Implementation Details on Matching Module
We implement our matching module described in Section 4.3 using the multi-head attention mechanism Vaswani et al. (2017b). To incorporate denoising timestep $t$, we adopt adaptive normalization Perez et al. (2018), modulating the output of the matching module using the denoising timestep. Specifically, we use the embedding of the query condition $g_{\tau}(y_{\tau})$, the support conditions $\{g_{\tau}(y^{i}_{\tau})\}_{i=1}^{N}$, and the support images $\{f(x^{i})\}_{i=1}^{N}$ as the query, key, and value inputs, respectively. Let $Q\in\mathbb{R}^{M\times d}$, $K,V\in\mathbb{R}^{(N\cdot M)\times d}$, and the timestep embedding be denoted as $t_{\text{emb}}\in\mathbb{R}^{d_{t}}$, our matching mechanism operates as below:
@@ A.2 Meta-training Settings
  • Checkpoint: We use the Stable Diffusion v1.5 checkpoint available from the HuggingFace CompVis (2022).
  • Hyperparameters: We train using the AdamW optimizer Loshchilov and Hutter (2019) with a learning rate of $1\times 10^{-5}$ and a weight decay of 0.01. For each training batch, we randomly select two tasks per batch, each accompanied by a support set of three example pairs sampled for its query condition.
  • Spatial condition representation: Following ControlNet Zhang et al. (2023), we represent all conditioning inputs as RGB images with a resolution of 512×512.
@@ A.3 Few-shot Fine-tuning Settings
When fine-tuning our meta-trained model on 30 support examples of the novel condition type, at most 600 fine-tuning steps with a batch size of 10 (approximately 1 hour on an RTX 3090 GPU) suffice for all six tasks. Early-stopping with a held-out validation example to monitor the denoising can be applied to avoid overfitting. We found that low-level conditions (e.g., edges) tend to converge faster than high-level conditions (e.g., pose). We also adopt AdamW optimizer Loshchilov and Hutter (2019) with a learning rate of $1\times 10^{-5}$ and a weight decay of 0.01.
@@ A.5 Dataset Construction
  • Training data: We randomly sample a subset of data from the LAION dataset Schuhmann et al. (2021). To identify images containing humans, we use the YOLO11x model Jocher et al. (2023). Based on this filtering, we construct a dataset consisting of 150K images with humans and 150K images without humans.
  • Evaluation data: For the Canny, HED, Depth, and Normal tasks, we evaluate our model on 5,000 images from the COCO2017 validation set Lin et al. (2014). For the Pose task, evaluation is limited to images where humans are successfully detected by the Openpose model Cao et al. (2017). Similarly, for the Densepose task, we only evaluate our model with images where the Densepose model Güler et al. (2018) detects a human.
@@ Appendix B Comparison with Training-Free Methods
We propose a few-shot framework for adapting to new spatial conditions in T2I diffusion models. A natural question arises: why use a few-shot approach when training-free methods exist Bansal et al. (2023); Yu et al. (2023); Mo et al. (2024); Lin et al. (2024). To illustrate the advantages of our method (UNet backbone), as mentioned earlier in Section 5.1, we compare it with Ctrl-X Lin et al. (2024) and FreeControl Mo et al. (2024) state-of-the-art training-free approaches capable of handling diverse spatial conditions without relying on additional pretrained networks.
To perform spatial control using a condition image, Ctrl-X Lin et al. (2024) replaces the features and attention maps when processing the noisy latent with the ones encoded from the noisy condition image. We re-implement Ctrl-X using the Stable Diffusion v1.5 backbone for a comparable setting. FreeControl Mo et al. (2024) controls image structure by constructing a PCA basis from object features, projecting both the condition and noisy image feature maps onto this basis, and minimizing an energy function that encourages alignment between the two projections during generation.
@@
We follow FreeControl’s original evaluation protocol and assess controllability of UFC and two training-free baselines on 30 images from the ImageNet-R-TI2I dataset Tumanyan et al. (2023), which includes 10 object categories with three captions each. FreeControl has limited flexibility in generating diverse object categories, as it requires constructing a separate PCA basis for each. This makes large-scale evaluation on 5,000 images from the COCO 2017 validation set Lin et al. (2014) impractical, since each prompt must be individually inspected to identify objects and build corresponding PCA bases.
As the evaluation dataset of FreeControl excludes human subjects, we compare UFC (30-shot) against FreeControl on four tasks: Canny, HED, Depth, and Normal. Due to the limited number of images, we do not report FID Heusel et al. (2017) or Inception Score Salimans et al. (2016), as they would not provide reliable estimates of image quality.
@@
Table 5: Controllability scores and average generation time of training-free baselines and UFC (Ours, 30-shot) on generating 30 images from ImageNet-R-TI2I Tumanyan et al. (2023).
@@
Figure 5: Qualitative comparison between UFC (30-shot) and FreeControl Mo et al. (2024), Ctrl-X Lin et al. (2024) on four control tasks.
@@ Appendix D Experiments with DiT Backbone
As discussed in Section 4.3, our framework can work with both UNet and DiT backbones. We extend UFC to the DiT architecture Peebles and Xie (2023), using Stable Diffusion v3 Esser et al. (2024) as an example, initialized from the publicly available v3.5-medium checkpoint AI (2024) on Hugging Face. This model consists of 24 Transformer layers in the diffusion backbone. We initialize the image encoder with pretrained weights of the diffusion backbone and keep it frozen during training. For the condition encoder, we initialize it as a trainable copy of the first 12 layers of the diffusion backbone. We train our model on 8 NVIDIA RTX A6000 GPUs.
@@
We follow the same optimizer settings, training dataset, and evaluation protocol described in Section  5.1. For inference, we use the flow-matching Euler scheduler introduced in Stable Diffusion 3Esser et al. (2024), with a classifier-free guidance (CFG) scale of 5.0, 28 generation steps, and a fixed seed of 42.
@@ Appendix E More Results
As previously discussed in Section 5.3, we evaluate UFC on more novel spatial conditions, including 3D meshes, wireframes, and point clouds. We use iso3d Ebert (2025), a 3D isolated object dataset(with no background), to generate spatial conditions for image pairs. We only report qualitative results, as the absence of a pre-trained condition-prediction network prevents us from measuring controllability using the evaluation approach in Section 5.1. Moreover, the insufficient collected validation sets hinder reliable assessment of image quality with FID Heusel et al. (2017) or Inception Score Salimans et al. (2016). Figure 11 shows the qualitative results of our method, demonstrating the effectiveness of our method in a few-shot (30 supports) setting.

### v2

@@ 5.2 Main Results
Fully-supervised ControlNet [58] 0.3598 0.5972 89.09 14.14 0.525 0.4824
Uni-ControlNet [59] 0.3378 0.5808 92.60 14.59 0.340 0.4695
Training-free Ctrl-X [29] 0.2901 0.3002 98.10 19.38 0.005 0.1352
Few-shot Prompt Diffusion [53] 0.2120 0.2887 98.81 20.34 0.012 0.2266
Uni-ControlNet + FT [59] 0.2222 0.2917 99.21 20.36 0.010 0.2687
Prompt Diffusion + FT [53] 0.2773 0.4810 95.64 18.04 0.034 0.3548
@@
Fully-supervised ControlNet [58] 21.06 17.15 20.16 21.73 44.33 34.72
Uni-ControlNet [59] 16.79 16.76 20.35 19.59 47.13 36.77
Training-free Ctrl-X [53] 29.83 30.18 30.32 30.35 56.60 45.88
Few-shot Prompt Diffusion [53] 24.85 23.00 24.07 23.28 43.60 34.15
Uni-ControlNet [59] + FT 24.58 25.48 24.77 25.12 47.82 36.06
Prompt Diffusion [53] + FT 20.57 19.86 24.56 24.30 45.94 36.83
@@
We report the 30-shot performance of our model with UNet backbone [41] in terms of two aspects: controllability (Table 1) and image quality (Table 2). Despite using only a few annotated examples, UFC achieves comparable image quality and over 90% of the controllability performance of fully supervised baselines across most conditions. We observe that dense condition maps benefit more from the matching mechanism, as fine-grained similarity modeling enables more effective control features. This is reflected in the smaller relative performance gap with fully supervised baselines on Densepose compared to Pose.
@@ 5.3 Analysis
To assess the effectiveness of our design in handling the large discrepancy across condition modalities, we conduct an ablation study with two variants of UFC on UNet backbone [41]: (1) UFC w/o Matching and (2) UFC w/o Fine-tuning. Both variants are trained on the same dataset as our main method and evaluated on the COCO 2017 validation split. In (1), we remove the matching mechanism and encode the query condition directly using the condition encoder i.e., $\mathcal{G}_{\tau}(y_{\tau})$ while retaining task-specific parameters and meta-learning paradigm. At test time, we fine-tune $\theta_{\tau}$ and $\mathcal{Z}$ using the small support set. In (2), we disable adaptation entirely by sharing all parameters across tasks during training and evaluating directly on unseen tasks without any fine-tuning.
@@
Since UFC is applicable to various diffusion backbones, we employ the stronger pre-trained model based on DiT [11] and compare it with the one based on UNet used in our main table. The quantitative results are reported in Table 4, showing that the DiT-based model with a stronger diffusion backbone consistently outperforms the UNet-based model in controllability across all tasks. The lower image quality compared to the UNet variant, which is also observed in [48], can be due to the mismatch between the optimized image resolution of the pre-trained diffusion backbone with the testing resolution (1024 and 512). Figure 9 in Appendix D shows that UFC (DiT) qualitatively achieves more fine-grained controllability than the UNet counterpart. More qualitative examples using the DiT backbone are shown in Figure 10 in Appendix.
@@
To further validate the few-shot capability of our method in more challenging settings, we evaluate our method on novel spatial conditions on 3D structures, such as 3D meshes, wireframes, and point clouds extracted from iso3d dataset [10]. Importantly, these control tasks involve not only novel spatial conditions but also different output image distributions from meta-training datasets, simulating a more realistic generalization scenario to unseen control tasks.
@@ A.1 Implementation Details on Matching Module
We implement our matching module described in Section 4.3 using the multi-head attention mechanism [51]. To incorporate denoising timestep $t$, we adopt adaptive normalization [37], modulating the output of the matching module using the denoising timestep. Specifically, we use the embedding of the query condition $g_{\tau}(y_{\tau})$, the support conditions $\{g_{\tau}(y^{i}_{\tau})\}_{i=1}^{N}$, and the support images $\{f(x^{i})\}_{i=1}^{N}$ as the query, key, and value inputs, respectively. Let $Q\in\mathbb{R}^{M\times d}$, $K,V\in\mathbb{R}^{(N\cdot M)\times d}$, and the timestep embedding be denoted as $t_{\text{emb}}\in\mathbb{R}^{d_{t}}$, our matching mechanism operates as below:
@@ A.2 Meta-training Settings
  • Checkpoint: We use the Stable Diffusion v1.5 checkpoint available from the HuggingFace [9].
  • Hyperparameters: We train using the AdamW optimizer [33] with a learning rate of $1\times 10^{-5}$ and a weight decay of 0.01. For each training batch, we randomly select two tasks per batch, each accompanied by a support set of three example pairs sampled for its query condition.
  • Spatial condition representation: Following ControlNet [58], we represent all conditioning inputs as RGB images with a resolution of 512×512.
@@ A.3 Few-shot Fine-tuning Settings
When fine-tuning our meta-trained model on 30 support examples of the novel condition type, at most 600 fine-tuning steps with a batch size of 10 (approximately 1 hour on an RTX 3090 GPU) suffice for all six tasks. Early-stopping with a held-out validation example to monitor the denoising can be applied to avoid overfitting. We found that low-level conditions (e.g., edges) tend to converge faster than high-level conditions (e.g., pose). We also adopt AdamW optimizer [33] with a learning rate of $1\times 10^{-5}$ and a weight decay of 0.01.
@@ A.5 Dataset Construction
  • Training data: We randomly sample a subset of data from the LAION dataset [45]. To identify images containing humans, we use the YOLO11x model [21]. Based on this filtering, we construct a dataset consisting of 150K images with humans and 150K images without humans.
  • Evaluation data: For the Canny, HED, Depth, and Normal tasks, we evaluate our model on 5,000 images from the COCO2017 validation set [30]. For the Pose task, evaluation is limited to images where humans are successfully detected by the Openpose model [6]. Similarly, for the Densepose task, we only evaluate our model with images where the Densepose model [14] detects a human.
@@ Appendix B Comparison with Training-Free Methods
We propose a few-shot framework for adapting to new spatial conditions in T2I diffusion models. A natural question arises: why use a few-shot approach when training-free methods exist [2, 57, 34, 29]. To illustrate the advantages of our method (UNet backbone), as mentioned earlier in Section 5.1, we compare it with Ctrl-X [29] and FreeControl [34] state-of-the-art training-free approaches capable of handling diverse spatial conditions without relying on additional pretrained networks.
To perform spatial control using a condition image, Ctrl-X [29] replaces the features and attention maps when processing the noisy latent with the ones encoded from the noisy condition image. We re-implement Ctrl-X using the Stable Diffusion v1.5 backbone for a comparable setting. FreeControl [34] controls image structure by constructing a PCA basis from object features, projecting both the condition and noisy image feature maps onto this basis, and minimizing an energy function that encourages alignment between the two projections during generation.
@@
We follow FreeControl’s original evaluation protocol and assess controllability of UFC and two training-free baselines on 30 images from the ImageNet-R-TI2I dataset [49], which includes 10 object categories with three captions each. FreeControl has limited flexibility in generating diverse object categories, as it requires constructing a separate PCA basis for each. This makes large-scale evaluation on 5,000 images from the COCO 2017 validation set [30] impractical, since each prompt must be individually inspected to identify objects and build corresponding PCA bases.
As the evaluation dataset of FreeControl excludes human subjects, we compare UFC (30-shot) against FreeControl on four tasks: Canny, HED, Depth, and Normal. Due to the limited number of images, we do not report FID [15] or Inception Score [44], as they would not provide reliable estimates of image quality.
@@
Table 5: Controllability scores and average generation time of training-free baselines and UFC (Ours, 30-shot) on generating 30 images from ImageNet-R-TI2I [49].
@@
Figure 5: Qualitative comparison between UFC (30-shot) and FreeControl [34], Ctrl-X [29] on four control tasks.
@@ Appendix D Experiments with DiT Backbone
As discussed in Section 4.3, our framework can work with both UNet and DiT backbones. We extend UFC to the DiT architecture [36], using Stable Diffusion v3 [11] as an example, initialized from the publicly available v3.5-medium checkpoint [1] on Hugging Face. This model consists of 24 Transformer layers in the diffusion backbone. We initialize the image encoder with pretrained weights of the diffusion backbone and keep it frozen during training. For the condition encoder, we initialize it as a trainable copy of the first 12 layers of the diffusion backbone. We train our model on 8 NVIDIA RTX A6000 GPUs.
@@
We follow the same optimizer settings, training dataset, and evaluation protocol described in Section  5.1. For inference, we use the flow-matching Euler scheduler introduced in Stable Diffusion 3[11], with a classifier-free guidance (CFG) scale of 5.0, 28 generation steps, and a fixed seed of 42.
@@ Appendix E More Results
As previously discussed in Section 5.3, we evaluate UFC on more novel spatial conditions, including 3D meshes, wireframes, and point clouds. We use iso3d [10], a 3D isolated object dataset(with no background), to generate spatial conditions for image pairs. We only report qualitative results, as the absence of a pre-trained condition-prediction network prevents us from measuring controllability using the evaluation approach in Section 5.1. Moreover, the insufficient collected validation sets hinder reliable assessment of image quality with FID [15] or Inception Score [44]. Figure 11 shows the qualitative results of our method, demonstrating the effectiveness of our method in a few-shot (30 supports) setting.